OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? Answer 3x − 3y = 11; 6y = 9x − 22 5y = -9x − 6; -27x + 15y = 15 4x − 7y = 7; -14y = 8x − 14 -4x + 2 = 3y; 12y − 8 = -16x please help me

OpenStudy (anonymous):

-4x + 2 = 3y; 12y − 8 = -16x

OpenStudy (jiteshmeghwal9):

the equation with\[\frac{a_1}{a_2}=\frac{b_1}{b_2}=\frac{c_1}{c_2}\]will have infinite solution